1) Is the following function linear? Can you evaluate the function for the given value of x.
f(x)=1 - x^2 ; f(0)
2)Can you help me write an equation of the line that passes through the given point and satisfies the given condition for the foolowing two problems:
a)(1,3) ; parallele to y=2x-1
b) (3,2) ; perpendicular to y= -3x
3)What is the property that the statement 4(2 + 3)=4(2) + 4(3) illustrates?

Answer by stanbon(29603)

1) Is the following function linear?
Ans: No, it is a quadratic equation.
--------------
Can you evaluate the function for the given value of x.
f(x)=1 - x^2 ;
Ans: f(0) = 1-0^2 = 1
----------------------------
2)Can you help me write an equation of the line that passes through the given point and satisfies the given condition for the foolowing two problems:
a)(1,3) ; parallel to y=2x-1
Equation form: y=mx+b; you need m and b
m=2 because the line you want is parallel to y=2x-1 which has slope = 2
Solve for b: 3 = 2*1+b ; b = 1
EQUATIOn:
y = 2x+1
------------------
b) (3,2) ; perpendicular to y= -3x
m = 1/3
2=(1/3)3+b ; b = 1
EQUATION:
y = (1/3)x + 1
--------------------
3)What is the property that the statement 4(2 + 3)=4(2) + 4(3) illustrates?
The distributive property.
====================
Cheers,
Stan H.